OpenStudy (anonymous):

Solve the equation for x: 3x + 6y = 15. Answer x = 15 – 6y x = 5 – 2y x = 12 – 6y x = 5 – 6y

OpenStudy (anonymous):

: 3x + 6y = 15. subtract 6y 3x=15-6y divide by 3 x=5-2y

OpenStudy (anonymous):

3x + 6y = 15 Move all terms containing x to the left, all other terms to the right. Add '-6y' to each side of the equation. 3x + 6y + -6y = 15 + -6y Combine like terms: 6y + -6y = 0 3x + 0 = 15 + -6y 3x = 15 + -6y Divide each side by '3'. x = 5 + -2y x = 5 + -2y

OpenStudy (anonymous):

x = 5 -2y